Z to the 2nd power - z = 12
Find the smaller solution of this equation.

Answer by checkley77(12844)

Z to the 2nd power - z = 12
z^2-z-12=0
(z-4)(z+3)=0
z-4=0
z=4 answer.
z+3=0
z=-3 answer.